# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

### v1

@@ Abstract
Decision trees are traditionally trained using greedy heuristics that locally optimize an impurity or information metric. Recently there has been a surge of interest in optimal decision tree (ODT) methods that globally optimize accuracy directly. We identify two relatively unexplored aspects of ODTs: the objective function used in training trees and tuning techniques. Additionally, the value of optimal methods is not well understood yet, as the literature provides conflicting results, with some demonstrating superior out-of-sample performance of ODTs over greedy approaches, while others show the exact opposite. In this paper, we address these three questions: what objective to optimize in ODTs; how to tune ODTs; and how do optimal and greedy methods compare? Our experimental evaluation examines 13 objective functions, including four novel objectives resulting from our analysis, seven tuning methods, and six claims from the literature on optimal and greedy methods on 165 real and synthetic data sets. Through our analysis, both conceptually and experimentally, we discover new non-concave objectives, highlight the importance of proper tuning, support and refute several claims from the literature, and provide clear recommendations for researchers and practitioners on the usage of greedy and optimal methods, and code for future comparisons.
@@ 1 Introduction
Decision trees (DTs) are among the most-used (interpretable) machine learning (ML) models. Despite their simplicity, they can learn complex non-linear relationships in data. Their human comprehensibility answers the need for interpretable models in high-stake domains (Rudin, 2019; Arrieta et al., 2020), provided the trees are small. Optimal decision trees (ODTs) specifically, which provably optimize an objective for a given size limit, provide small but accurate models on many tabular datasets and thus combine high performance with interpretability (Piltaver et al., 2016; Loh, 2014; Carrizosa et al., 2021).
Because training optimal decision trees with respect to a size limit is NP-Hard (Hyafil and Rivest, 1976), most early decision tree learning methods were greedy top-down induction heuristics, such as CART (Breiman et al., 1984) and C4.5 (Quinlan, 1993), which locally optimize some impurity or information gain metric for each branching node.
Consequently, greedy decision tree learning has been extensively studied given its long history; important examples being splitting criteria (Mingers, 1989b; Buntine and Niblett, 1992; Shih, 1999; Raileanu and Stoffel, 2004; Wang and Xia, 2017) and pruning techniques to avoid overfitting (Mingers, 1989a; Patil et al., 2010; Esposito et al., 1997).
In contrast, optimal decision tree research is a much younger field, with the last decade seeing major advancements. The main topic of research has been improving scalability by reducing runtimes and supporting larger datasets. Researchers have employed a variety of techniques such as (mixed) integer programming (MIP) (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017), constraint programming (Verhaeghe et al., 2020), satisfiability (Hu et al., 2020; Narodytska et al., 2018) and dynamic programming with bounds (Aglin et al., 2020a; Lin et al., 2020; Demirović et al., 2022). Whereas ten years ago optimal methods were limited to small data sets, due to algorithmic advancements and an increase in computation power, we can now use a recent dynamic programming approach (van der Linden et al., 2023) to analyze data sets with up to hundreds of thousands of samples.
@@
This motivated us to conduct a thorough review of existing decision tree methods, both greedy and optimal, focusing on objective functions used during training and different tuning approaches. Based on our findings, we motivate new objective functions and tuning approaches that are specific to optimal methods that globally optimize the objective. To support our study, we conducted the largest evaluation to this date concerning optimal and greedy decision tree methods, taking into account 13 different objective functions, seven tuning approaches, 165 real-world and synthetic data sets (small and large), and trees that go beyond small tree-depth limits. This provides us with a wealth of data to analyze and insights into the strengths and weaknesses of greedy and optimal approaches.
@@
  • In Section 3, we analyze and compare nine existing greedy decision tree accuracy objectives for optimal decision trees. We observe that the strict concavity of many of these objectives, as required by greedy top-down inducting approaches (Kearns and Mansour, 1996), yields counter-intuitive results when trained to optimality. Since ODTs do not require concave objectives, we contribute four new non-concave objectives. Our experiments show that these new non-concave objectives outperform the default accuracy objective for smaller data sets by more effectively penalizing leaf nodes with nearly uniform class distributions. For larger data sets, ODTs can directly optimize the in-sample accuracy.
  • In Section 4, we compare seven complexity tuning approaches for ODTs, four of which were proposed before, and three new tuning approaches. Our experiments highlight the importance of tuning optimal decision tree methods, and that (surprisingly) the differences between the commonly used tuning methods are small.
  • In Section 5, we review previous comparisons between greedy and optimal approaches, formulate best practices for future comparisons, and provide data and code to support proper benchmarking. [^1] The code will be made public on paper publication. We apply these practices in evaluating six claims from the literature on the performance of greedy and optimal trees. Specifically, we evaluate the following claims:
@@ 2. Claim 2:
      Optimal methods have a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
@@ 6. Claim 6:
      The expected depth of greedy trees remains (in practice) close to the optimal tree depth, even for complex (large) true tree size (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
    Our results support Claims 126, and refute Claims 34 and 5.
@@ 2 Related Work
This section introduces previous literature on decision tree learning with a focus on optimal methods. For longer reviews, we refer to surveys by Safavian and Landgrebe (1991); Kotsiantis (2013) and Costa and Pedreira (2023).
Decision tree learning started several decades ago with AID (Morgan and Sonquist, 1963), a recursive approach to regression analysis, later adapted for classification in CHAID (Kass, 1980). Since then, two of the most popular decision tree learning algorithms have been CART (Breiman et al., 1984) and ID3 (Quinlan, 1986) with its successors C4.5 (Quinlan, 1993). Each of these uses top-down induction (TDI) to greedily partition the data by finding a split that is locally optimal according to an information or impurity criterion. Overfitting is prevented either by early stopping rules such as a minimum information gain or by post-pruning the tree. We discuss splitting criteria in Section 3.1.
@@
Because computing trees optimally is NP-Hard (Hyafil and Rivest, 1976; Cox et al., 1989; Murphy and McCraw, 1991), historically most approaches have been heuristics. Although a few early optimal dynamic programming (Schumacher and Sevcik, 1976; Payne and Meisel, 1977; Miyakawa, 1985; Cox et al., 1989; Nijssen and Fromont, 2007, 2010) and an extreme point tabu search (Bennett and Blue, 1996) approaches have been proposed, only recently, with increasing computation power and new algorithmic improvements, interest in computing optimal trees resurged.
This resurge started with mixed-integer programming (MIP) formulations for ODTs (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017) with several consecutive improvements (Verwer and Zhang, 2019; Aghaei et al., 2021; Liu et al., 2024; Günlük et al., 2021; Zhu et al., 2020; Hua et al., 2022; Alès et al., 2024). The advantages of these MIP methods are that they can find splits with arbitrary thresholds on the continuous features and can easily be adapted by adding linear constraints or changing the objective, including objectives that operate on the whole tree instead of summing the objectives of independent leaf nodes. The disadvantage is poor scalability because of a weak linear relaxation and the inability of the MIP solver to recognize the independence between subtrees.
@@
Similar techniques are exploited in dynamic-programming (DP) based approaches, which a recent survey (Costa and Pedreira, 2023) indicates as the most promising approach in terms of scalability. DP exploits the independent subtree structure and reuses partial solutions to repeated subproblems. The addition of bound-based pruning (Aglin et al., 2020a, b), improved lower bounds (Hu et al., 2019; Lin et al., 2020; Demirović et al., 2022; Chaouki et al., 2024), and a faster subprocedure for trees of depth two (Demirović et al., 2022) have greatly improved the scalability of the basic DP approach. The advantages of DP are the good scalability for realistic use cases, specifically with respect to the number of instances. The disadvantages are the need for binarization and an exponential runtime with respect to the number of features and maximum tree size.
Recent developments for ODTs are incorporating continuous features (Mazumder et al., 2022; Shati et al., 2023); exploring the Rashomon set of all close to optimal models (Xin et al., 2022; Semenova et al., 2023); quantifying and reducing explanation redundancy (Izza et al., 2022; Audemard et al., 2022); improving anytime performance (Kiossou et al., 2022; Demirović et al., 2023); improving memory usage (Aglin et al., 2022); and applying ODTs to other objectives such as regression (Zhang et al., 2023; van den Bos et al., 2024), quantile regression (Lemaire et al., 2024), fairness constraints (Aghaei et al., 2019; Jo et al., 2022; van der Linden et al., 2022), robustness (Vos and Verwer, 2022; Justin et al., 2022), survival analysis (Zhang et al., 2024; Huisman et al., 2024), prescriptive policy generation (Bertsimas et al., 2019; Jo et al., 2021; van der Linden et al., 2023), and learning MDP policies (Vos and Verwer, 2023).
@@ 3 The Optimization Objective for Optimal Decision Trees
Decision tree learning objectives typically optimize two parts: some accuracy objective (i.e., accuracy or one of its proxies, such as information gain) and a tree-complexity objective (e.g., number of nodes). Some objectives only include the first, while others, such as the minimum description length (see below), combine both into one. In this section, we focus on the first: the accuracy objective. In Section 4, we discuss the tree-complexity objective.
@@
Section 3.1 lists existing greedy splitting criteria and how they relate to ODT objectives. In our analysis of these objectives in Section 3.2, we observe that the concavity traditionally required by greedy heuristics (Kearns and Mansour, 1996) is not helpful when training ODTs. Therefore, we introduce four novel non-concave objectives in Section 3.3. We then empirically compare all these objectives in Section 3.4 and discuss our findings in Section 3.5. Our main finding is that optimal and greedy decision tree learners benefit from optimizing different objectives. Greedy learners perform best with concave objectives while optimal learners achieve the best performance when optimizing non-concave objectives.
3.1 From Greedy Criteria to Objectives
Greedy TDI criteria have been extensively analyzed and compared. We extend this comparison to optimal decision trees. In the following, we first consider how existing greedy criteria relate to objective functions for ODTs.
Greedy top-down induction (TDI) methods typically consider two types of criteria: a local splitting criterion that decides how to split a node and a pruning criterion that decides which nodes to keep during pruning. Although the goal of decision tree learning is to maximize accuracy, TDI methods rarely directly optimize accuracy but instead optimize a proxy. One of the first classification tree methods, CHAID (Kass, 1980), maximizes the reduction in the $\chi^{2}$ statistic; CART (Breiman et al., 1984) optimizes the Gini impurity; and ID3 (Quinlan, 1986) and C4.5 (Quinlan, 1993) optimize information gain (or entropy). Some greedy methods use the same criterion for splitting and pruning; others employ another criterion for pruning such as the minimum description length (Mehta et al., 1995) or the pessimistic error rate in C4.5 (Quinlan, 1993). We explain each in more detail below.
ODTs, on the other hand, do not have a splitting or pruning criterion but optimize a global objective (also called a loss function), typically expressed as a sum over the errors of the leaf nodes plus some term that penalizes the size of the tree. Almost all ODT methods directly optimize the accuracy (Nijssen and Fromont, 2007; Bertsimas and Dunn, 2017; Verwer and Zhang, 2017, 2019; Hu et al., 2019; Aglin et al., 2020a; Hu et al., 2020; Aghaei et al., 2021). To improve the performance on unbalanced data sets, some have considered other objectives such as balanced accuracy (Lin et al., 2020) or F1-score (Demirović and Stuckey, 2021). Nijssen and Fromont (2010) have implemented the pessimistic error objective from C4.5 and a Bayesian objective function, but do not discuss the effect on the out-of-sample performance when these objectives are used.
We now present a list of splitting and pruning criteria used for TDI heuristics and for each provide an equivalent formulation as an ODT leaf node objective. We rewrite these functions as a function $f$ with as input the leaf node size $n$ and the number of misclassifications $e$. Let $L$ be a list of leaf nodes of size $n$ and with $e$ misclassifications, then the minimization objective of the whole tree becomes
$\sum_{(n,e)\in L}f(n,e)\,.$
Accuracy:
The most direct way to optimize out-of-sample accuracy is to optimize the in-sample accuracy by minimizing the misclassification score. The leaf node objective in terms of the leaf node size $n$ and the number of misclassifications $e$ is
$f(n,e)=e\,.$
A possible disadvantage of this objective is its ‘flatness’: it cannot distinguish between many different solutions. For example, $f(4,2)+f(6,0)=f(5,1)+f(5,1)$.
Gini impurity:
Commonly used in greedy top-down heuristics is the Gini impurity. Gini impurities are from 0 to 1. Weighted Gini impurity scores are obtained by multiplying the Gini impurity by the number of instances in that leaf node. Let $p_{0}=\frac{e}{n}$ denote the probability of the first class and $p_{1}=\frac{n-e}{n}$ the probability of the second class. Then the objective value is
$f(n,e)=n(1-p_{0}^{2}-p_{1}^{2})\,.$
Square root Gini:
$f(n,e)=n\sqrt{1-p_{0}^{2}-p_{1}^{2}}\,.$
Entropy:
The second commonly used objective in top-down heuristics is entropy (or information gain). Like the Gini impurity, weighted scores are obtained by multiplying by the size of the leaf node:
$f(n,e)=-\frac{n}{2}(p_{0}\log_{2}p_{0}+p_{1}\log_{2}p_{1})\,.$
Minimum error:
Niblett (1987) estimates the expected error for nodes by assuming that every class has equal probability. With $n$ samples in a node and $|\mathcal{K}|$ the number of labels, and $n_{c}$ the count of the majority label, the expected error rate is
$\frac{n-n_{c}+|\mathcal{K}|-1}{n+|\mathcal{K}|}\,.$
Since we consider binary classification, $|\mathcal{K}|=2$, and $n_{c}=n-e$. Therefore:
$f(n,e)=\frac{n(e+1)}{n+2}$
Pessimistic error:
$f(n,e)=e+\frac{1}{2}\,.$
Binomial pessimistic error (Binom.):
Minimum description length (Quinlan):
$L(n,e,b)=\ln{(b+1)}+\ln{\binom{n}{e}}\,.$ (1)
$f(n,e)=L(n,e,\lfloor(n+1)/2\rfloor)\,.$
Minimum description length (Mehta):
Bayesian:
$f(n,e)=\frac{B(e+\rho_{0},n-e+\rho_{1})}{B(\rho_{0},\rho_{1})}\,.$
3.2 Analysis
@@
The reason why TDI methods do not optimize accuracy directly is that an accuracy splitting criterion is often unable to find an improving split in unbalanced data. When splitting a node, TDI methods evaluate all possible splits and choose the split that minimizes the splitting criterion value for the resulting class distributions among the new nodes of each possible split. Fig. 1 visually explains why using accuracy as a splitting criterion is worse at distinguishing improving splits than Gini impurity or entropy. The left side shows the function values for accuracy, Gini impurity, and entropy. The right side shows how a locally optimal split can be found geometrically. When splitting a node with a probability of the first class of $p$ into two new nodes with probabilities $p_{1}$ and $p_{2}$, the new weighted splitting criterion value can be found by drawing a straight line from the criterion value at point $p_{1}$ to $p_{2}$. The intersection of the straight line at $p$ is the sum of the weighted criterion value of the two nodes. For Gini impurity and entropy, this value is always lower than the criterion value of the parent node, because both functions are strictly concave. Accuracy, however, is not strictly concave, and when $p<0.5,p_{1}<0.5$, and $p_{2}<0.5$ (or equivalently all are greater than $0.5$), the weighted sum of the criterion value of the child nodes is the same as that of the parent node. Moreover, for any values $p_{1}<0.5$ and $p_{2}<0.5$ the weighted sum of the criterion values is the same, and therefore no distinction can be made between these splits. Thus TDI heuristics require strictly concave splitting criteria (Kearns and Mansour, 1996) and therefore do not optimize accuracy directly.
@@
Figure 2: Objective values for different objective functions for a single leaf node. Left, the leaf node size is fixed at $n=100$. Right, the misclassifications are fixed at $e=20$. Surprisingly, the value of the strictly concave objectives increases for a fixed error and increasing leaf node size.

### v2

@@ Abstract
Recently there has been a surge of interest in optimal decision tree (ODT) methods that globally optimize accuracy directly, in contrast to traditional approaches that locally optimize an impurity or information metric. However, the value of optimal methods is not well understood yet, as the literature provides conflicting results, with some demonstrating superior out-of-sample performance of ODTs over greedy approaches, while others show the opposite. Through a novel extensive experimental study, we provide new insights into the design and behavior of learning decision trees. In particular, we identify and analyze two relatively unexplored aspects of ODTs: the objective function used in training trees, and tuning techniques. Thus, we address these three questions: what objective to optimize in ODTs; how to tune ODTs; and how do optimal and greedy methods compare? Our experimental evaluation examines 11 objective functions, six tuning methods, and six claims from the literature on optimal and greedy methods on 180 real and synthetic data sets. Through our analysis, both conceptually and experimentally, we show the effect of (non-)concave objectives in greedy and optimal approaches; we highlight the importance of proper tuning of ODTs; support and refute several claims from the literature; provide clear recommendations for researchers and practitioners on the usage of greedy and optimal methods; and code for future comparisons.
@@ 1 Introduction
Decision trees (DTs) are among the most-used (interpretable) machine learning (ML) models. Despite their simplicity, they can learn complex non-linear relationships in data and their human comprehensibility answers the need for interpretable models in high-stake domains (Rudin, 2019; Arrieta et al., 2020), provided the trees are small. Optimal decision trees (ODTs) specifically, which provably optimize an objective for a given size limit, provide small but accurate models on many tabular data sets and thus combine high performance with interpretability (Piltaver et al., 2016; Loh, 2014; Carrizosa et al., 2021).
Because training optimal decision trees with respect to a size limit is NP-hard (Hyafil and Rivest, 1976), most early decision tree learning methods were greedy top-down induction heuristics. Such methods, like CART (Breiman et al., 1984) and C4.5 (Quinlan, 1993), locally optimize some impurity or information gain metric for each branching node.
Consequently, greedy decision tree learning has been extensively studied given its long history; important examples being splitting criteria (Mingers, 1989b; Buntine and Niblett, 1992; Shih, 1999; Raileanu and Stoffel, 2004; Wang and Xia, 2017) and pruning techniques to avoid overfitting (Mingers, 1989a; Esposito et al., 1997; Patil et al., 2010).
In contrast, optimal decision tree research is a much younger field, with the last decade seeing major advancements. The main topic of research has been improving scalability by reducing runtimes and supporting larger data sets. Researchers have employed a variety of techniques such as (mixed) integer programming (MIP) (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017), constraint programming (Verhaeghe et al., 2020), Boolean satisfiability (Hu et al., 2020; Narodytska et al., 2018) and dynamic programming with bounds (Aglin et al., 2020a; Lin et al., 2020; Demirović et al., 2022). Whereas ten years ago optimal methods were limited to small data sets, due to algorithmic advancements and an increase in computation power, we can now use a recent dynamic programming approach (van der Linden et al., 2023) to analyze data sets with up to hundreds of thousands of samples.
@@
This motivated us to conduct a thorough experimental evaluation of existing decision tree methods, both greedy and optimal, focusing on objective functions used during training and different tuning approaches. Based on our findings, we motivate new objective functions and tuning approaches that are specific to optimal methods that globally optimize the objective. To support our study, we conducted the largest evaluation to this date concerning optimal and greedy decision tree methods, taking into account 11 different objective functions, six tuning approaches, 180 real-world and synthetic data sets (small and large), and trees that go beyond small tree-depth limits. This provides us with a wealth of data to analyze and improves our understanding of how to apply greedy and optimal approaches for training decision trees.
@@
  • In Section 3, we analyze and experimentally compare nine existing greedy decision tree accuracy objectives. Since we observe that the strict concavity of these objectives, as required by greedy top-down inducting approaches (Kearns and Mansour, 1996), is counterproductive when trained to optimality, we also introduce and experiment with two new non-concave objectives. Our experiments show that greedy and optimal approaches respond oppositely to the (non-)concavity of these objectives. Additionally, we show the benefit of objectives that include a regularizing component for noisy data (in addition to the regular tree size tuning).
  • In Section 4, we compare six complexity tuning approaches for ODTs, four of which were proposed before, and two new tuning approaches that we introduce here. Our experiments highlight the importance of tuning optimal decision tree methods, and that (surprisingly) the accuracy differences between the commonly used tuning methods are small, although there are differences in resulting tree size and runtimes.
  • In Section 5, we analyze previous comparisons between greedy and optimal approaches, formulate best practices for future comparisons, and provide data and code to support proper benchmarking. [^1] The code will be made public on paper publication. We apply these practices in evaluating six claims from the literature on the performance of greedy and optimal trees:
@@ 2. Claim 2:
      Optimal methods obtain a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
@@ 6. Claim 6:
      The question length of greedy trees remains (in practice) close to that of optimal trees (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
    Our results support Claims 12, and 6, and refute Claims 34, and 5.
@@ 2 Related Work
This section introduces previous literature on decision tree learning with a focus on optimal methods. For longer reviews, we refer to surveys by Safavian and Landgrebe (1991); Kotsiantis (2013); Costa and Pedreira (2023) and Blockeel et al. (2023).
Decision tree learning started several decades ago with AID (Morgan and Sonquist, 1963), a recursive approach to regression analysis, later adapted for classification in CHAID (Kass, 1980). Since then, two of the most popular decision tree learning algorithms have been CART (Breiman et al., 1984) and ID3 (Quinlan, 1986) with its successor C4.5 (Quinlan, 1993). Each of these uses top-down induction (TDI) to greedily partition the data by finding a split that is locally optimal according to an information or impurity criterion. Overfitting is prevented either by early stopping rules such as a minimum information gain or by post-pruning the tree. We discuss splitting criteria in Section 3.1.
@@
Because computing optimal trees is NP-hard (Hyafil and Rivest, 1976; Cox et al., 1989; Murphy and McCraw, 1991), historically most approaches have been heuristics. Although a few early optimal dynamic programming (Schumacher and Sevcik, 1976; Payne and Meisel, 1977; Miyakawa, 1985; Cox et al., 1989; Nijssen and Fromont, 2007, 2010) and an extreme-point tabu search (Bennett and Blue, 1996) approaches were proposed, only recently, with increased compute and algorithmic advancements, interest in optimal trees resurged.
This resurge started with mixed-integer programming (MIP) formulations for ODTs (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017) with several consecutive improvements (Verwer and Zhang, 2019; Zhu et al., 2020; Günlük et al., 2021; Hua et al., 2022; Alès et al., 2024; Aghaei et al., 2024; Liu et al., 2024). The advantages of these MIP methods are that they can find splits with arbitrary thresholds on the continuous features and can easily be adapted by adding linear constraints or changing the objective, including objectives that operate on the whole tree instead of summing the objectives of independent leaf nodes. The disadvantage is poor scalability because of a weak linear relaxation and the inability of the MIP solver to recognize the independence between subtrees.
@@
Similar techniques are exploited in dynamic-programming (DP) based approaches, which a recent survey (Costa and Pedreira, 2023) indicates as the most promising approach in terms of scalability. DP exploits the independent subtree structure and reuses partial solutions to repeated subproblems. The addition of bound-based pruning (Aglin et al., 2020a, b), improved lower bounds (Hu et al., 2019; Lin et al., 2020; Demirović et al., 2022), and a faster subprocedure for trees of depth two (Demirović et al., 2022) have greatly improved the scalability of the basic DP approach. The advantages of DP are the good scalability for realistic use cases, specifically with respect to the number of instances. The disadvantages are the need for binarization and an exponential runtime with respect to the number of features and maximum tree size.
Recent developments for ODTs are incorporating continuous features (Mazumder et al., 2022; Shati et al., 2023; Brita et al., 2025); exploring the Rashomon set of all close to optimal models (Xin et al., 2022; Semenova et al., 2023); quantifying and reducing explanation redundancy (Izza et al., 2022; Audemard et al., 2022); improving anytime performance (Kiossou et al., 2022; Demirović et al., 2023); improving memory usage (Aglin et al., 2022); and applying ODTs to other objectives such as regression (Zhang et al., 2023; van den Bos et al., 2024), quantile regression (Lemaire et al., 2024), fairness constraints (Aghaei et al., 2019; van der Linden et al., 2022; Jo et al., 2023), robustness (Vos and Verwer, 2022; Justin et al., 2022), survival analysis (Zhang et al., 2024; Huisman et al., 2024), prescriptive policy generation (Bertsimas et al., 2019; Jo et al., 2021; van der Linden et al., 2023), and learning MDP policies (Vos and Verwer, 2023).
Alternatively, others have studied decision trees with soft (probabilistic) decision splits, also known as randomized trees (Blanquero et al., 2021, 2022), or trees with oblique (multivariate) splits (Bertsimas and Dunn, 2017; Zhu et al., 2020; Blanquero et al., 2021; Boutilier et al., 2023; Engür and Soylu, 2024). However, such models are less human-comprehensible and out-of-scope for this paper.
@@ 3 The Optimization Objective for Optimal Decision Trees
Decision tree learning objectives typically optimize two parts: some accuracy objective (i.e., accuracy or one of its proxies, such as information gain) and a tree-complexity objective (e.g., number of nodes). In this section, we focus on the first: the accuracy objective. In Section 4, we discuss the tree-complexity objective.
@@
Section 3.1 explains why existing greedy splitting criteria do not optimize accuracy directly. We transform these greedy criteria to ODT objectives and observe in our analysis of these objectives in Section 3.2, that the strict concavity traditionally required by greedy heuristics (Kearns and Mansour, 1996) is not helpful when training ODTs. Therefore, we introduce two novel non-concave objectives in Section 3.3. We then empirically compare all these objectives in Section 3.4 and discuss our findings in Section 3.5. Our main finding is that greedy learners perform best with strictly concave splitting criteria, whereas optimal learners achieve the best performance with non-strictly-concave objectives.
3.1 Greedy Proxies for Accuracy
Although the goal of decision tree learning is to maximize accuracy, TDI methods rarely optimize accuracy directly but instead optimize a proxy, such as the reduction in the $\chi^{2}$-statistic in CHAID (Kass, 1980), Gini impurity by CART (Breiman et al., 1984), and information gain (entropy) by ID3 (Quinlan, 1986) and C4.5 (Quinlan, 1993).
@@ 3.1 Greedy Proxies for Accuracy
The reason why TDI methods do not optimize accuracy directly is that an accuracy splitting criterion is often unable to find an improving split in unbalanced data. When splitting a node, TDI methods evaluate all possible splits and choose the split that minimizes the splitting criterion value for the resulting class distributions among the new nodes of each possible split. Fig. 1 visually explains why using accuracy as a splitting criterion is worse at distinguishing improving splits than Gini impurity or entropy. The left side shows the function values for accuracy, Gini impurity, and entropy for binary classification. The right side shows how a locally optimal split can be found geometrically. When splitting a node with a probability of the first class of $p$ into two new nodes with probabilities $p_{1}$ and $p_{2}$, the new weighted splitting criterion value can be found by drawing a straight line from the criterion value at point $p_{1}$ to $p_{2}$. The intersection of the straight line at $p$ is the sum of the weighted criterion value of the two nodes. For Gini impurity and entropy, this value is always lower than the criterion value of the parent node, because both functions are strictly concave. Accuracy, however, is not strictly concave, and when $p\leq 0.5,p_{1}\leq 0.5$, and $p_{2}\leq 0.5$ (or equivalently, all are greater than or equal to $0.5$), the weighted sum of the criterion value of the child nodes is the same as that of the parent node. Moreover, for any values $p_{1}\leq 0.5$ and $p_{2}\leq 0.5$ the weighted sum of the criterion values is the same, and therefore no distinction can be made between these splits. Thus TDI heuristics require strictly concave splitting criteria (Kearns and Mansour, 1996) and therefore do not optimize accuracy directly.
@@
Figure 2: Objective values for different objective functions for a single leaf node. (Left) The leaf node size is fixed at $n=100$. (Right) The misclassifications are fixed at $e=20$. Surprisingly, the value of the strictly concave objectives increases for a fixed error and increasing leaf node size.
3.2 Analysis of ODT Objectives
To increase our understanding of greedy and optimal decision tree learning approaches, we analyze accuracy and eight other existing greedy splitting and pruning criteria and rewrite them as ODT objectives: Gini impurity, square root Gini, entropy, minimum error, binomial pessimistic error (Binom.), minimum description length (MDL, two encodings: Quinlan and Mehta), and Bayesian. In Appendix A, we rewrite each of these as a function $f$ with as input the number of instances $n$ that reach this leaf and the number of misclassifications $e$ in this leaf. Let $L$ be a set of leaf nodes of size $n$ and with $e$ misclassifications, then the minimization objective of the whole tree becomes $\sum_{(n,e)\in L}f(n,e)$.